# Technologies for Modulation of Visible and Near-UV Light and their Applications

### v1

@@ I.3 Biology
Visible wavelength optical systems play a critical role in the spatiotemporal control and readout of cells and their dynamics within complex biological systems. Light can be used to target optical transitions of molecules and proteins within cells to trigger cell functions or read out cell states [105]. These techniques typically rely on microscopes and table-top optical setups including liquid crystal spatial light modulators, galvanometer mirrors, digital micromirrors (Section II.1), and acousto-optic deflectors (Section II.2.1) to generate and reconfigure three-dimensional optical patterns. Due to the size, complexity, and speed of traditional free-space optical systems, biological experiments are often restricted to isolated tissues or cells or in-vivo studies of immobilized animals. High-speed visible photonic integrated circuits offer a path towards miniaturization and scalability of these optical systems enabling implantable, wearable, and portable devices. Here we outline two biological applications that require high-speed, low power, and scalable visible modulation: optogenetic neural stimulation and high resolution microscopy.
@@ I.3.1 Optogenetic Neural Stimulation
These actuators and reporters rely on optical transitions that lie in the visible wavelength range from 450 nm to 600 nm [105]. For neural activity control, an opsin is used to render a neuron light-sensitive for optogenetic stimulation: when a neuron is exposed to light within the opsin’s excitation bandwidth, an action potential is either excited or inhibited. Channelrhodopsin-2 (ChR2), a common opsin, is typically switched between 1-10 Hz, while ChETA, a high-speed opsin, can switch up to 200 Hz. Although the speed requirement for individual neuron control is slow, high-speed excitation (i.e. MHz or GHz) could enable efficient, large-volume excitation. The amount of optical power necessary for one photon optogenetic excitation is approximately 5-10 $\mu$W per cell.
@@
The scaling of active implantable photonic devices is ultimately limited by power consumption, drive voltages, and speed of the modulators which, in turn, control the complexity of optical patterns that can be produced. The optical pattern complexity directly impacts the number of single neurons or neuron populations that can be addressed. Although biological processes generally occur at millisecond timescales (Hz), the faster speeds (MHz-GHz) allow for addressing larger populations and real-time, closed-loop interrogation of the biological systems as biological processes and behaviors are complex and dynamic. It will be necessary to limit power consumption and drive voltages of active devices in proximity to biological systems to prevent disruption to natural processes or obfuscation of small neural signals. Although these circuits may remain outside of the brain, depending on dimensions of the probe and heat sinking methods, residual increases in temperatures along the shank can occur as the number of modulators increases. An ultimate limit is imposed by the brain temperature which cannot be raised more than ${}^{\circ}$ C. This demands that much less than a mW of power is dissipated per modulator for a truly large-scale system. The future potential for wireless or battery operation of photonic devices will further restrict the power consumption and drive voltages needed for reconfiguration.
@@ I.3.2 High-resolution Microscopy
High-resolution microscopy techniques rely on 3D patterning and structuring of light to overcome limitations in resolution, FOV, and photo-damage during biological imaging, revealing sub-cellular structure and dynamics. Fluorescence microscopy uses fluorescent-labeled proteins or the endogenous label-free fluorescence within a sample. These can span wavelengths from UV through the visible spectrum. For example, light-sheet and structured illumination fluorescence microscopy both rely on shaped excitation to improve performance. In light-sheet microscopy, a Gaussian or Bessel beam is formed orthogonal to the detection path to provide a sheet of localized fluorescence signal across a wide FOV, leading to faster, larger volumetric imaging with high axial resolution and reduced photodamage by eliminating out-of-focus signal [128, 129, 130]. In structured illumination microscopy, sinusoidal excitation patterns are used to form translated or rotated Moire fringes in which high frequency information within the sample can be revealed leading to super-resolution imaging [131]. However, the typical table-top apparatus to generate and scan the excitation light is bulky, limiting the volumes and types of samples that can be studied. The miniaturization of these techniques could provide a path for high-resolution microscopy to become more widely accessible and eventually used for diagnostics, point-of-care settings, or even wearable devices for continuous monitoring.
@@
Recently, thermally-controlled aluminum oxide photonic integrated circuits have been used for far-field UV label-free structured illumination microscopy [135, 136]. The phase-shifters are used to precisely control the phases of multiple beams for high contrast interference above the chip. Super-resolution is achieved using a standard wide-field microscope in conjunction with the chip illumination. Although the power constraints for imaging may not be as restrictive as implantable devices, similar power and voltage drive constraints remain to prevent disruption of biological processes. The reconfiguration speeds of excitation light will impact imaging frame rates and volumes, so high-speed visible photonic integrated circuits that push beyond the speeds of state-of-the-art acousto-optic deflectors and spatial light modulators remains a critical goal.
@@ I.4 Summary
All three of the applications discussed here rely on active control of UV and visible light but each specific application has different performance requirements. To ground our discussion of the current state-of-the-art in visible light modulation, we focus on five metrics of interest: (1) modulation speed; (2) control power per channel; (3) size per channel; (4) optical bandwidth; and (5) technological maturity. Fig. 2(a) shows a qualitative description of the importance of each of these metrics for the three application areas and Fig. 2(b) shows the wavelengths of interest. Ideally one technology would provide high speed modulation with low power across the full UV to visible range in a compact form factor. However, there are often trade-offs between these parameters as will become clear in Sec. II and Sec. III. It is likely that each application will require a bespoke or hybrid solution with optimized system-level performance.
@@ II Dynamic Surfaces and Bulk Modulation
Here, we discuss bulk devices and reflective surfaces which enable full control of the intensity and phase distribution of free-space beams. These devices often can control the spatial distribution of the output light as well as the intensity, phase, and frequency of individual beams. They are currently mainstays in the application areas discussed above and provide excellent modulation metrics, though are bulky and often require high power. Still, they provide inspiration for the integrated devices discussed in the next section and there is ongoing work to improve the optical specifications while reducing the size and power consumption.
@@ II.1 Dynamic Reflective Surfaces
In a dynamic reflective surface, an array of controllable reflective pixels modifies the characteristics of an incoming beam of light. These systems are mature and commercially available.
@@ II.1.1 Mirror Arrays
Mirrors are a fundamental optical element and can modify the direction and phase of an incident laser beam. The orientation of the mirror controls the position of the reflected beam and the position in the travel direction controls the phase. By actuating the stage on which a mirror is mounted, the position and phase of the incoming laser light is thus modulated. A simple example is the mirror galvanometer which is used for rapidly scanning a single laser beam or image. These are advanced commercial technologies used across industries including confocal microscopy and laser beam machining. These systems are simple, fairly fast, and reliable. However, they are bulky, imprecise, and usually limited to a single or few channels. Alternatively, the mirror can be mounted on a piezo-acutated stage enabling positioning and pointing with nanometer accuracy. This is an indispensable tool for high-resolution feedback control and optical interferometers including LIGO [137].
The advances of micro-electromechanical systems (MEMS) technology has enabled miniaturization of the actuators and the electronic control. MEMS-controlled mirrors reduce required voltages and increase the modulation bandwidth due to their micron-scale size. With these advances, MEMS-controlled mirrors can approach MHz modulation speeds. Such MEMS mirrors can be integrated into a single-mode fiber for binary on/off control or can be assembled into compact one or two-dimensional arrays to form programmable surfaces with locally reconfigurable reflection properties controlled by applying pixel-by-pixel voltages. See [138] for a full review of both commercial and academic work on micromirror arrays. Combined with free-space propagation, optical Fourier transforms, and additional spatial filtering these reflective surfaces can be used to control the full image plane [139] and are thus referred to as spatial light modulators (SLMs).
MEMS-mirror-based SLMs can control the amplitude or phase of the reflected light. For amplitude control, the orientation of a MEMS-mirror is rotated in a digital or analog fashion. The digital control selectively directs a micron-scale portion of the incident beam to the target image position or to a beam dump, producing the desired intensity pattern. These amplitude-only SLMs are called digital micro-mirror devices (DMDs). The DMDs were initially used for digital light projection (DLP) technology as licensed by Texas Instruments (TI), but have widely found applications in complex wavefront shaping for optical microscopy [140, 141, 142, 143], lithography [144], trapped-ion control [55], optical tweezer generation [145], and high-precision atomic potential generation [146, 147] based on their high-extinction ratio between on/off orientations. These DMDs can provide high temporal (32 kHz) and spatial (4K UHD) resolution but the modulation is generally inefficient as it relies on removing a portion of incoming light to change the far-field amplitude distribution.
@@
Another important class of mechanical systems is deformable mirrors (DMs). Thousands of MEMS actuators deform the reflective surfaces introducing a local phase on the reflected wavefront. Unlike the DMDs or GLVs used as amplitude SLMs, DMs modulate the phase of the reflected beam, which is lossless in principle. Thus, the main applications of DMs are in optical imaging for example to actively compensate for biological sample-induced aberrations in optical microscopy [149] or distortion due to an inhomogeneous atmosphere [150]. Here, the number of actuators determines the degree of control. The modulation speed is comparable to DMDs but the number of actuators is limited to a few thousands.
@@ II.1.2 Liquid Crystal Arrays
When a phase SLM with high spatial resolutions is required, a surface composed of nematic liquid crystal on silicon (LCoS) can be used at the cost of slower frame rate of tens to hundreds of hertz. The refractive index of the LC medium depends on the orientation of the LC molecules with respect to the propagation direction of light. LC molecules align to an external electric field and thus the orientation can be tuned by applying a voltage across the LC region. As the LC molecules rotate, the refractive index of the LC medium changes, imparting a voltage-dependent phase delay on the reflected beam. These devices have been extensively reviewed previously [151, 152]. We discuss integrated versions of these devices in Sec. III.4.
@@ II.2 Transmissive Modulators
II.2.1 Acoustic-Optic Modulators
The acoustic-optic modulator (AOM) is the workhorse for many applications, including communication, Q-switching in lasers, and fast control of atomic systems for quantum technologies. In an AOM, Brillouin scattering [153] is used to modulate the position, amplitude, frequency, and phase of an incident laser [154]. A piezoelectric transducer generates an acoustic wave in a crystal from an input radiofrequency (rf) signal. This acoustic wave creates a periodic modulation in the refractive index of the crystal, acting as a diffraction grating for an incident laser beam. The diffracted component of the laser inherits the phase, frequency, and amplitude of the rf signal and the position is determined by the modulation frequency.
This frequency-dependent positioning can be used to easily separate modulated and un-modulated portions of the beam. It can also be used to provide spatial control. An AOM can be used as an imaging device by carefully engineering the rf frequency components [155]. Moreover, the acousto-optic modulation based on slow shear waves enables deflecting the incident beam into a wide range of angles with hundreds of resolvable spots [156]. These acousto-optic deflectors are central for neutral-atom quantum information processors as discussed in Sec. I.
Current commercial systems boast diffraction efficiencies of 1000:1 and switching speeds of below 10 ns. However, due to their bulk nature, watts of rf power is needed per channel [157] which is unfavorable for independent control of a large number of laser beams. As discussed further in Sec. III.3, there has been some initial work to integrate these systems to reduce the size and power and hybrid solutions are discussed in Sec. IV.
@@ II.2.2 Electro-optic Modulators
Electro-optic (EO) modulation broadly refers to a refractive index change due to an external electric field though the exact physical mechanism varies and depends on the specific material [158]. Many materials exhibit electro-optic nonlinearities and fast EO modulation is routine at telecom wavelengths. In particular, Lithium Niobate (LiN) has long been the workhorse for bulk light modulation. LiN has a relatively large bandgap (3.8 eV), promising low-loss propagation down to around 350 nm. Additionally, bulk LiN boasts a strong Pockels effect ($r_{33}\sim 31$ pm/V)  [159] which can be used for EO modulation. Conventional LiN modulators use either bulk LiN or large waveguides defined by titanium diffusion or annealed proton exchange in bulk LiN. Due to the low index contrast and thus confinement, these devices are bulky and require high voltages. Nevertheless they out-compete current Si-based modulators, especially for applications with no power or size constraints. In contrast to acoustic-optic modulation, EO modulation is not used to control the spatial degree of freedom and the modulated and unmodulated light cannot easily be separated.
@@ II.3 Summary and Outlook
Figure 3: Qualitative depiction of the capabilities of acoustic optic modulators, liquid crystal spatial light modulators, and micromirror spatial light modulators for each of the five figures of merit discussed in this review.
All the applications discussed in Sec. I rely on the dynamic surfaces and bulk modulation technologies discussed in this section. These systems are mature with well-engineered commercial solutions that enable quick integration into larger application systems. However, they are still quite bulky. Although, acoustic modulation can provide high speed single-channel modulation, the large per-channel size and rf power are prohibitive. In contrast, dynamic surfaces provide full image control with a single device but are generally limited to tens of kHz. Fig. 3 shows a qualitative depiction of the three technologies discussed - micromirror dynamic surfaces, liquid crystal dynamic surfaces, and acoustic optic modulators - in the context of the figures of merit of this paper - technological maturity, modulation speed, electrical control power, size, and optical bandwidth. Fig. 3 puts the optical bandwidth and speed of these commercial technologies in the context of all the technologies discussed in this review.
While these bulk technologies provide an easy way to modulate the amplitude, phase, and often spatial distribution of light, they clearly lack the integrability in terms of size and power consumption needed for next generation quantum, biological, and augmented-reality applications. As the optical pixel-size is reduced the electronic control may become the ultimate resolution bottleneck.

### v2

@@ I.3 Biology
Visible wavelength optical systems play a critical role in the spatiotemporal control and readout of cells and their dynamics within complex biological systems. Light can be used to target optical transitions of molecules and proteins within cells to trigger cell functions or read out cell states [108]. These techniques typically rely on microscopes and table-top optical setups including liquid crystal spatial light modulators, galvanometer mirrors, digital micromirrors (Section II.1), and acousto-optic deflectors (Section II.2.1) to generate and reconfigure three-dimensional optical patterns. Due to the size, complexity, and speed of traditional free-space optical systems, biological experiments are often restricted to isolated tissues or cells or in-vivo studies of immobilized animals. High-speed visible photonic integrated circuits offer a path towards miniaturization and scalability of these optical systems enabling portable, wearable, and implantable devices. Here we outline two biological applications that require high-speed, low power, and scalable visible modulation: optogenetic neural stimulation and high resolution microscopy. Schematics of these two applications using active intgrated photonic circuits for light delivery is shown in Fig. 3.
Figure 3: a) Implantable optogenetic neural probe based on a photonic chip for control of neural activity through multicolor excitation. b) photonic chip based far-field structured illumination for high resolution imaging.
@@ I.3.1 Optogenetic Neural Stimulation
These actuators and reporters rely on optical transitions that lie in the visible wavelength range from 450 nm to 600 nm [108] as shown in Fig. 4. For neural activity control, an opsin is used to render a neuron light-sensitive for optogenetic stimulation: when a neuron is exposed to light within the opsin’s excitation bandwidth, an action potential is either excited or inhibited. Channelrhodopsin-2 (ChR2), a common opsin, is typically switched between 1-10 Hz, while ChETA, a high-speed opsin, can switch up to 200 Hz. Although the speed requirement for individual neuron control is slow, high-speed excitation (i.e. MHz or GHz) could enable efficient, large-volume excitation. The amount of optical power necessary for one photon optogenetic excitation is approximately 5-10 $\mu$W per cell.
@@
The scaling of active implantable photonic devices is ultimately limited by power consumption, drive voltages, and speed of the modulators which, in turn, control the complexity of optical patterns that can be produced. The optical pattern complexity directly impacts the number of single neurons or neuron populations that can be addressed. Although biological processes generally occur at millisecond timescales (Hz), the faster speeds (MHz-GHz) allow for addressing larger populations and real-time, closed-loop interrogation of the biological systems as biological processes and behaviors are complex and dynamic. It will be necessary to limit power consumption and drive voltages of active devices in proximity to biological systems to prevent disruption to natural processes or obfuscation of small neural signals. Although these circuits may remain outside of the brain, depending on dimensions of the probe and heat sinking methods, residual increases in temperatures along the shank can occur as the number of modulators increases. An ultimate limit is imposed by the brain temperature which cannot be raised more than 1-2 C. This demands that much less than a mW of power is dissipated per modulator for a truly large-scale system. The future potential for wireless or battery operation of photonic devices will further restrict the power consumption and drive voltages needed for reconfiguration.
@@ I.3.2 High-resolution Microscopy
High-resolution microscopy techniques rely on 3D patterning and structuring of light to overcome limitations in resolution, FOV, and photo-damage during biological imaging, revealing sub-cellular structure and dynamics. Fluorescence microscopy uses fluorescent-labeled proteins or the endogenous label-free fluorescence within a sample. These can span wavelengths from UV through the visible spectrum. For example, light-sheet and structured illumination fluorescence microscopy both rely on shaped excitation to improve performance. In light-sheet microscopy, a Gaussian or Bessel beam is formed orthogonal to the detection path to provide a sheet of localized fluorescence signal across a wide FOV, leading to faster, larger volumetric imaging with high axial resolution and reduced photodamage by eliminating out-of-focus signal [131, 132, 133]. In structured illumination microscopy, sinusoidal excitation patterns are used to form translated or rotated Moire fringes in which high frequency information within the sample can be revealed leading to super-resolution imaging [134]. However, the typical table-top apparatus to generate and scan the excitation light is bulky, limiting the volumes and types of samples that can be studied. The miniaturization of these techniques could provide a path for widely accessible high-resolution microscopy, eventually used for diagnostics, point-of-care settings, or even wearable devices for continuous monitoring.
@@
Recently, thermally-controlled aluminum oxide photonic integrated circuits have been used for far-field UV label-free structured illumination microscopy [138, 139]. The phase-shifters are used to precisely control the phases of multiple beams for high contrast interference above the chip. Super-resolution is achieved using a standard wide-field microscope in conjunction with chip illumination. Although the power constraints for imaging may not be as restrictive as implantable devices, similar power and voltage drive constraints remain to prevent disruption of biological processes. The reconfiguration speed of the excitation light will impact imaging frame rates and volumes, so high-speed visible photonic integrated circuits that push beyond the speeds of state-of-the-art acousto-optic deflectors and spatial light modulators remains a critical goal.
@@ I.4 Summary
All three of the applications discussed here rely on active control of visible light but each specific application has different performance requirements. To ground our discussion of the current state-of-the-art in visible light modulation, we focus on five metrics of interest: (1) modulation speed; (2) control power per channel; (3) size per channel; (4) optical bandwidth; and (5) technological maturity. Fig. 4(a) shows a qualitative description of the importance of each of these metrics for the three application areas and Fig. 4(b) shows the wavelengths of interest for applications of visible light modulation. Ideally one technology would provide high speed modulation with low power across the full UV to visible range in a compact form factor. However, there are often trade-offs between these parameters as will become clear in Sec. II and Sec. III. It is likely that each application will require a bespoke or hybrid solution with optimized system-level performance.
@@ II Dynamic Surfaces and Bulk Modulation
Here we discuss reflective surfaces and bulk devices which enable full control of the intensity and phase distribution of free-space beams. These devices often can control the spatial distribution of the output light as well as the intensity, phase, and frequency of individual beams. They are currently mainstays in the application areas discussed above and provide excellent modulation metrics, though are bulky and often require high power, and there is ongoing work to improve the optical specifications while reducing the size and power consumption.
@@ II.1 Dynamic Reflective Surfaces
In a dynamic reflective surface, an array of controllable reflective pixels modifies the characteristics of an incoming beam of light. Combined with free-space propagation, optical Fourier transforms, and additional spatial filtering these reflective surfaces can be used to control the full image plane [140] and are thus generally referred to as spatial light modulators (SLMs). As discussed below, the modulation mechanism can be chosen to optimize for speed, optical power efficiency, and degree of control. These systems are mature and commercially available. Spatial light modulators have been used in a plethora of applications including direct-write lithography [141], optical communication [142, 143], fiber amplifiers [144], and microscopy [145]. Research is ongoing to improve pixel density and speed.
@@ II.1.1 Mirror Arrays
Mirrors are a fundamental optical element and can modify the direction and phase of an incident laser beam: the orientation of the mirror controls the position of the reflected beam and the position in the travel direction controls the phase. Actuating the stage on which a mirror is mounted thus modulates the position and phase of the incoming laser light. A simple example is the mirror galvanometer which is used for rapidly scanning a single laser beam or image. These are advanced commercial technologies used across industries including confocal microscopy and laser beam machining and are simple, fairly fast, and reliable. However, they are bulky, imprecise, and usually limited to a single or few channels. Alternatively, the mirror can be mounted on a piezo-acutated stage enabling positioning and pointing with nanometer accuracy. This is an indispensable tool for high-resolution feedback control and optical interferometers, including LIGO [146].
The advances of micro-electromechanical systems (MEMS) technology has enabled miniaturization of the actuators and the electronic control. MEMS-controlled mirrors reduce required voltages and increase the modulation bandwidth due to their micron-scale size. With these advances, MEMS-controlled mirrors can approach MHz modulation speeds. Such MEMS mirrors can be integrated into a single-mode fiber for binary on/off control or can be assembled into compact one or two-dimensional arrays to form programmable surfaces with locally reconfigurable reflection properties controlled by applying pixel-by-pixel voltages. See [147] for a full review of both commercial and academic work on micromirror arrays.
MEMS-mirror-based SLMs can control the amplitude or phase of the reflected light. For amplitude control, the orientation of a MEMS-mirror is rotated in a digital or analog fashion. The digital control selectively directs a micron-scale portion of the incident beam to the target image position or to a beam dump, producing the desired intensity pattern. These amplitude-only SLMs are called digital micromirror devices (DMDs). The DMDs were initially used for digital light projection (DLP) technology as licensed by Texas Instruments (TI), but have widely found applications in complex wavefront shaping for optical microscopy [148, 149, 150, 151], lithography [152], trapped-ion control [57], optical tweezer generation [153], and high-precision atomic potential generation [154, 155]. DMDs provide high temporal (32 kHz) and spatial (4K UHD) resolution and high-extinction ratio between on/off orientations. However, the modulation is generally inefficient as it relies on removing a portion of the incoming light to change the far-field amplitude distribution.
@@
Another important class of mechanical systems is deformable mirrors (DMs). Thousands of MEMS actuators deform the reflective surfaces introducing a local phase on the reflected wavefront. Unlike the DMDs or GLVs used as amplitude SLMs, DMs modulate the phase of the reflected beam, which is lossless in principle. Thus, the main applications of DMs are in optical imaging to actively compensate aberrations, for instance for due to biological samples in optical microscopy [157] or distortion due to an inhomogeneous atmosphere [158]. Here, the number of actuators determines the degree of control. The modulation speed is comparable to DMDs but the number of actuators is limited to a few thousands.
@@ II.1.2 Liquid Crystal Arrays
When a phase SLM with high spatial resolution is required, a surface composed of nematic liquid crystal on silicon (LCoS) can be used. This comes at the cost of slower frame rate of tens to hundreds of hertz. The refractive index of the LC medium depends on the orientation of the LC molecules with respect to the propagation direction of light. LC molecules align to an external electric field and thus the orientation can be tuned by applying a voltage across the LC region. As the LC molecules rotate, the refractive index of the LC medium changes, imparting a voltage-dependent phase delay on the reflected beam. These devices have been extensively reviewed previously [159, 160] and we discuss integrated versions of these devices in Sec. III.5.
@@ II.2 Transmissive Modulators
II.2.1 Acousto-Optic Modulators
The acousto-optic modulator (AOM) underpins many applications, including communication, Q-switching in lasers, and fast control of atomic systems for quantum technologies. In an AOM, Brillouin scattering [161] is used to modulate the position, amplitude, frequency, and phase of an incident laser [162]. A piezoelectric transducer generates an acoustic wave in a crystal from an input radiofrequency (rf) signal. This acoustic wave creates a periodic modulation in the refractive index of the crystal, acting as a controllable diffraction grating for an incident laser beam. The diffracted component of the laser inherits the phase, frequency, and amplitude of the rf signal and the displacement angle is determined by the modulation frequency.
The frequency-dependent positioning can be used to easily separate modulated and un-modulated portions of the beam. An AOM can be used as an imaging device by carefully engineering the rf frequency components [163]. Moreover, acousto-optic modulation based on slow shear waves enables deflecting the incident beam into a wide range of angles with hundreds of resolvable spots [164]. These acousto-optic deflectors are central for neutral-atom quantum information processors as discussed in Sec. I.
Current commercial systems boast diffraction efficiencies of 1000:1 and switching speeds of below 10 ns. However, due to their bulk nature, watts of rf power is needed per channel [165] which is unfavorable for independent control of a large number of laser beams. As discussed further in Sec. III.4, there has been some initial work to integrate these systems to reduce the size and power. Hybrid solutions are discussed in Sec. IV.
@@ II.2.2 Electro-Optic Modulators
Electro-optic (EO) modulation broadly refers to a refractive index change due to an external electric field though the exact physical mechanism varies and depends on the specific material [166]. Many materials exhibit electro-optic nonlinearities and fast EO modulation is routine at telecom wavelengths. In particular, Lithium Niobate (LiNbO3 ) has long been the workhorse for bulk light modulation. LiNbO3 has a relatively large bandgap (3.8 eV), promising low-loss propagation down to around 350 nm, and bulk LiNbO3 boasts a strong Pockels effect ($r_{33}\sim 31$ pm/V)  [167] which can be used for EO modulation. Conventional LiNbO3 modulators use either bulk LiNbO3 or large waveguides with a weak change in the refractive index defined by titanium diffusion or annealed proton exchange in bulk LiNbO3  [168]. Due to the low index contrast and thus confinement, these devices are bulky and require high voltages. Nevertheless they out-compete current Si-based modulators, especially for applications with no power or size constraints. In contrast to AO modulation, EO modulation can be much faster but does not affect the spatial degree of freedom so the modulated and unmodulated light cannot easily be separated.
@@ II.3 Summary and Outlook
Figure 5: Qualitative depiction of the capabilities of bulk acousto-optic and electro-optic modulators (AOM, EOM), and liquid crystal and micromirror spatial light modulators (LC SLM, MM SLM) for each of the five figures of merit discussed in this review.
All the applications discussed in Sec. I currently rely on the dynamic surfaces and bulk modulation technologies discussed in this section. These systems are mature with well-engineered commercial solutions that enable quick integration into larger application systems. However, no technology provides high speed, compact control. AOMs provide high speed single-channel modulation but the large per-channel size and rf power are prohibitive. In contrast, dynamic surfaces provide full image control with a single device but are generally limited to tens of kHz. Fig. 5 shows a qualitative depiction of the four technologies discussed - micromirror spatial light modulators (MM SLMs), liquid crystal spatial light modulators (LC SLMs), and bulk AOMs and EOMs - in the context of the figures of merit of this paper - technological maturity, modulation speed, electrical control power, size, and optical bandwidth. Fig. 9 puts the optical bandwidth and speed of these commercial technologies in the context of all the technologies discussed in this review.
While these bulk technologies provide an easy way to modulate the amplitude, phase, and often spatial distribution of light, they clearly lack the integrability in terms of size and power consumption needed for next generation quantum, biological, and augmented-reality applications. Current research is ongoing to provide full spatial-temporal control over light with the necessary speed, and optical bandwidth with reduced electrical power consumption. Optical phase arrays [169] are one solution to provide fast, spatial control of light. These rely on visible active integrated photonics which are discussed below in Sec. III. Active metasurfaces have also been demonstrated for spatial light modulation [170]. Finally, resonant surfaces can also improve speed – a recent work uses photonic crystal cavity pixels, achieving nanosecond pixel switching for a 2D array [171]. As the optical pixel-size is reduced electronic wiring and crosstalk may become the ultimate resolution bottleneck, though this could be overcome with further optical engineering [31].